OpenStudy (anonymous):

find the indicated outputs for (f)x=5x2-4x f(0)=

jimthompson5910 (jim_thompson5910):

f(x) = 5x^2-4x f(0) = 5(0)^2-4(0) f(0) = 5(0)-4(0) f(0) = 0-0 f(0) = 0